# RECAST: Strengthening LLMs’ Complex Instruction Following with Constraint-Verifiable Data

### v3

@@ Abstract
Large language models (LLMs) are increasingly expected to tackle complex tasks, driven by their expanding applications and users’ growing proficiency in crafting sophisticated prompts. However, as the number of explicitly stated requirements increases (particularly more than $10$ constraints), LLMs often struggle to accurately follow such complex instructions. To address this challenge, we propose RECAST, a novel framework for synthesizing datasets where each example incorporates far more constraints than those in existing benchmarks. These constraints are extracted from real-world prompt-response pairs to ensure practical relevance. RECAST enables automatic verification of constraint satisfaction via rule-based validators for quantitative constraints and LLM-based validators for qualitative ones. Using this framework, we construct RECAST-3030K, a large-scale, high-quality dataset comprising 3030k instances spanning $15$ constraint types. Experimental results demonstrate that models fine-tuned on RECAST-3030K show substantial improvements in following complex instructions. Moreover, the verifiability provided by RECAST enables the design of reward functions for reinforcement learning, which further boosts model performance on complex and challenging tasks.
@@
1 Introduction
Large language models (LLMs) have demonstrated remarkable capabilities in solving various NLP tasks Ahn et al. (2024); Chakrabarty et al. (2024); Gómez-Rodríguez and Williams (2023); Huang and Chang (2023); Jiang et al. (2024a); Zhao et al. (2023); Zheng et al. (2023) and are widely used in practical scenarios Gu et al. (2024); Liu et al. (2024a); Sun (2023); Thirunavukarasu et al. (2023). However, they still struggle with complex multi-constraint scenarios that are prevalent in real-world applications, such as constitutional AI systems requiring adherence to multiple principles simultaneously Bai et al. (2022); Findeis et al. (2024); Kundu et al. (2023); Kyrychenko et al. (2025), and enterprise assistants managing detailed business rules De Michele et al. (2025); Grohs et al. (2023); Robino (2025). As shown in Figure 1, when instruction complexity increases, with multiple explicit constraints in a single prompt, even advanced models like GPT-4o Achiam et al. (2023) show marked performance degradation, which has been emphasized in multiple benchmarks Ferraz et al. (2024); Jiang et al. (2024b); Qin et al. (2024); Wen et al. (2024).
While recent research efforts have focused on improving instruction-following capabilities through automated data generation Liang et al. (2024a); Liu et al. (2025, 2024b); Pham et al. (2024); Qi et al. (2024); Sun et al. (2024); Wang et al. (2023); Xu et al. (2024); Zhou et al. (2024) and improved training methodologiesCheng et al. (2024); Dong et al. (2024); He et al. (2024); Pham et al. (2024); Sun et al. (2024); Wallace et al. (2024); Wang et al. (2024a), several challenges remain underexplored. First, existing instruction datasets primarily consist of examples with simple constraints Pham et al. (2024); Wang et al. (2023); Xu et al. (2024) that inadequately reflect the richness and complexity of real-world queries. Second, conventional training methodologies prioritize overall response quality He et al. (2024); Pham et al. (2024) over fine-grained constraint adherence, leaving models prone to violating specific requirements when confronted with intricate, multi-constraint instructions.
To address the first challenge, we introduce RECAST (Realistic Extraction of Constraints for Augmented inStruction synThesis) that automatically generates complex instruction-following data endowed with multiple verifiable constraints, where each constraint is explicitly verifiable through either rule-based validators or LLM-based validators. To address the second challenge, we develop RLVC (Reinforcement Learning via Verifiable Constraints as rewards) to further enhance the model’s complex instruction-following capabilities, a reinforcement learning approach that provides fine-grained feedback on individual constraint satisfaction, enabling the model to optimize for multiple objectives simultaneously.
Based on RECAST, we constructed a high-quality, diverse, and constraint-verifiable dataset, RECAST-30K, which encompasses 15 common types of constraints. Using only 30K data, the base model, such as Qwen2.5-7B-Base Yang et al. (2024) trained on RECAST-30K outperforms the corresponding instruct model, which was trained on a larger-scale dataset. The RLVC approach further enhances constraint compliance, yielding substantial improvements in multi-constraint satisfaction. Through detailed analysis, we identify key factors affecting instruction-following reliability, finding that both constraint types and quantity impact model performance in complex scenarios. To summarize the highlights of our study:
  • We propose RECAST, a novel data-synthesis framework that constructs training examples infused with diverse, verifiable constraints.
  • We introduce RLVC, a reinforcement-learning approach that leverages constraint-specific reward signals for simultaneous multi-objective optimisation.
@@ 1 Introduction
2 Related Work
Complex Instruction Datasets Construction
Recent research has explored various approaches to automatically construct complex instruction datasets. WizardLM Xu et al. (2024) introduced Evol-Instruct, which iteratively increases instruction difficulty through rewriting. Conifer Sun et al. (2024) employs a top-down approach using GPT-4 to generate multi-level constraint instructions. ULTRAIF An et al. (2025) decomposes instructions into basic queries and constraints before recombination, while AIR Liu et al. (2025) applies an iterative refinement framework that progressively adds constraints through a model-judge cycle. Other approaches include MUFFIN Lou et al. (2023), which builds multi-faceted tasks by expanding input features, and CRaB Qi et al. (2024), which infers new constraints from existing instruction-response pairs through back-translation. Unlike these methods, our RECAST approach uniquely focuses on collecting diverse constraint types from high-quality responses and incorporates explicit verification mechanisms for both rule-based and model-based constraints, addressing limitations in existing methods regarding constraint complexity, diversity, and quality assurance.
Complex Instruction Following Capabilities Improvement
Recent algorithmic innovations have significantly advanced instruction following capabilities. AutoIF Dong et al. (2024) implements a self-dialogue generation process with execution-based verification, where models produce instructions, verification code, and test cases, filtering training data through executable feedback. RNR Wang et al. (2024a) extracts roles and rules from existing instructions to generate rule-compliant responses, while SPaR Cheng et al. (2024) employs self-play tree search refinement through multiple rounds of model self-adversarial interaction to optimize response quality. The approach in He et al. (2024) utilizes discriminative generation, applying a discriminator model to filter generated samples for higher-quality supervision. Unlike these methods, our RLVC framework specifically leverages constraint verifiability, designing multi-channel reward signals that treat each constraint as an independent optimization objective. This fine-grained reinforcement learning approach enables more precise feedback on constraint satisfaction, substantially improving both the accuracy and consistency of complex instruction following.
Figure 2: Overview of the RECAST framework and RLVC. Left: The RECAST pipeline generates complex instruction-following data through four steps: (1) seed data collection across diverse domains, (2) constraint construction with both rule-based and model-based verification methods, (3) instruction enhancement by integrating selected constraints, and (4) response synthesis ensuring constraints are satisfied. Top-right: Using RECAST-generated data to fine-tune LLMs through supervised learning. Bottom-right: RLVC framework leveraging constraint-specific verification to provide fine-grained rewards, guiding model optimization toward satisfying multiple constraints simultaneously.
@@ 3 Realistic Extraction of Constraints for Augmented inStruction synThesis
We present RECAST, a comprehensive framework for improving large language models’ ability to follow complex, multi-constraint instructions. As shown in Figure 2, the RECAST pipeline operates through a three-stage process. First, we construct a rich constraint pool by extracting both rule-based and model-based constraints from high-quality examples. Next, we synthesize enhanced instructions by selecting and naturally integrating appropriate constraints into original prompts. Finally, we generate consistent, constraint-compliant responses to these augmented instructions.
@@ 3.1 Seed Data Collection
We collect a diverse corpus of instruction-following examples spanning multiple domains, including code generation, creative writing, and factual question answering. This comprehensive dataset provides a rich foundation for identifying and extracting constraints that characterize high-quality responses across various task types.
@@ 3.2 Constraint Pool Construction
Rule-based Constraints Construction.
Rule-based constraints represent objective requirements that can be verified through deterministic methods, including structural elements (paragraph count), lexical specifications (keyword inclusion/exclusion), and quantitative parameters (word limits). These constraints are particularly valuable as they provide unambiguous verification signals for model training and evaluation. To systematically extract these constraints, we implement eight specialized rule-based extractors (detailed in Appendix D.1) that analyze responses to identify verifiable properties. These extractors recognize specific syntactic patterns, keyword frequencies, numerical parameters, and structural characteristics that can be programmatically verified. This programmatic approach ensures that each rule-based constraint has a corresponding verification method, providing clear training signals for models to learn specific response characteristics.
Model-based Constraints Construction.
Model-based constraints encompass subjective requirements that necessitate semantic understanding or qualitative judgment, such as stylistic elements (formality level), tonal qualities (politeness), and content characteristics (persuasiveness). These constraints are crucial for capturing the nuanced aspects of human communication that cannot be reduced to simple rules. Our construction process follows three key steps: First, we analyze each high-quality response to determine which constraint types from our taxonomy of 15 categories (detailed in Appendix D.2) are applicable, ensuring relevance to the specific content. Second, we employ LLMs to generate multiple concrete constraint instances for each applicable type, creating a comprehensive initial pool of model-based constraints. Third, we implement a filtering process to verify that each constraint is actually satisfied by its corresponding response, removing those that cannot be fulfilled. To validate the accuracy of filtering process, we conducted human evaluation on randomly selected constraint-response pairs, confirming that our methodology successfully identifies and retains only appropriate, verifiable constraints (detailed results in Appendix B.1).
@@ 3.3 Constraint-Augmented Instruction Synthesis
Constraint Selection.
We employ LLMs to select a coherent subset of constraints from the pool that are relevant and applicable to each original instruction. This selection process ensures that only contextually appropriate constraints are considered for integration, avoiding the inclusion of irrelevant or contradictory requirements.
Instruction Enhancement.
Once appropriate constraints are identified, we task multiple LLMs with integrating the selected constraints into the original instruction. Each model produces an enhanced instruction where constraints are incorporated naturally and coherently into the text.
Optimal Instruction Selection.
To ensure quality and consistency, we implement a majority voting mechanism wherein multiple LLMs evaluate and rank the candidate integrated instructions. The evaluation criteria include linguistic fluency, semantic coherence, and constraint completeness. We select the highest-ranked instruction as the final constraint-augmented instruction. We also performed random sampling on optimal instruction selection for human evaluation, the results of human evaluation are shown in Appendix B.2.
@@ 3.4 Instruction-Consistent Response Synthesis
To address potential inconsistencies between constraint-augmented instructions and original responses, we implement a two-phase process to ensure alignment while maintaining high response quality.
Diverse Response Generation.
We generate multiple candidate responses to each constraint-augmented instruction using distinct LLMs. This multi-model approach introduces diversity into the candidate pool and mitigates potential biases inherent to any single model. Each model generates responses independently, ensuring a broad representation of possible approaches to satisfying the complex instruction.
Response Quality Assessment.
We implement a robust majority voting mechanism to evaluate candidate responses based on multiple criteria, including linguistic fluency, factual accuracy, and comprehensive constraint adherence. This systematic evaluation process enables us to identify and select the response that most effectively fulfills all specified requirements while maintaining natural language quality. The winning response becomes part of the final dataset, paired with its corresponding enhanced instruction. To verify the quality of enhanced response, we conducted a human evaluation on randomly sampled instruction-response pairs. The detailed results of this evaluation are presented in Appendix B.3, confirming the high quality of our selected responses.
This methodical approach to generation and selection ensures that our final dataset comprises high-quality instruction-response pairs that demonstrate successful navigation of complex, multi-constraint scenarios. Each pair is annotated with its full set of constraints and corresponding validation method, facilitating downstream application and evaluation.
@@ 4 Reinforcement Learning with Verifiable Constraints as Rewards
Building on RECAST, we introduce RLVC, which leverages the verifiable nature of constraints to provide fine-grained reward signals during policy optimization. This approach enables more targeted feedback on constraint satisfaction, further improving models’ ability to simultaneously address multiple complex requirements.
@@ 4.1 Constraint Verification Mechanisms
We implement a dual-mode verification scheme that enables precise evaluation of constraint satisfaction. Each constraint in an instruction is assessed through either rule-based or LLM-based verification methods.
For rule-based constraints, we employ rule-based validators $V_{\text{rule-based}}(x,y,c_{i})$ that programmatically verify objective requirements through deterministic procedures. For model-based constraints that require evaluation of subjective qualities, we leverage LLM-based validators $V_{\text{model-based}}(x,y,c_{i})$. These validators assess more nuanced requirements that resist codification into explicit codes.
We formalize the constraint verification function for an instruction $x$ with response $y$ and constraint $c_{i}$ as:
@@ 4.2 Verifiable Constraints as Reward Signals
The key innovation of RLVC lies in its exploitation of RECAST’s constraint verifiability feature. Traditional reinforcement learning approaches typically rely on a single, holistic reward signal that fails to identify which specific constraints are violated. In contrast, RECAST-30K provides individually verifiable constraints, enabling us to design a more informative reward mechanism.
For an instruction $x$ containing multiple constraints $C=\{c_{1},c_{2},...,c_{n}\}$, we define the reward for a generated response $y$ as the average satisfaction rate across all constraints:
@@
This design provides an independent reward channel for each constraint, offering fine-grained feedback that guides the model toward simultaneously satisfying multiple constraints.
4.3 Policy Optimization
We employ Group Relative Policy Optimization (GRPO) Shao et al. (2024) as our policy learning algorithm. GRPO generates multiple candidate responses for the same instruction and computes advantage estimates through intra-group comparisons, allowing the model to learn the relative quality differences among various responses to the same instruction.
During policy updates, GRPO maximises the following objective:
where:

### v4

# RECAST: Expanding the Boundaries of LLMs’ Complex Instruction Following with Multi-Constraint Data

@@ Abstract
Large language models (LLMs) are increasingly expected to tackle complex tasks, driven by their expanding applications and users’ growing proficiency in crafting sophisticated prompts. However, as the number of explicitly stated requirements increases (particularly more than $10$ constraints), LLMs often struggle to accurately follow such complex instructions, which limits their applicability in complex real-world scenarios. To the best of our knowledge, existing datasets do not exceed 10 constraints per instance. To address this challenge, we propose RECAST, an efficient and scalable framework for synthesizing datasets where each example incorporates far more constraints than those in existing benchmarks, aiming to challenge and extend the boundaries of models’ ability to follow complex instructions. These constraints are extracted from real-world prompt-response pairs to ensure practical relevance. Using this framework, we construct RECAST-3030K, a large-scale, high-quality dataset comprising 3030k instances spanning $19$ constraint types. Experimental results demonstrate that models fine-tuned on RECAST-30K substantially improve in following complex instructions while maintaining their general capabilities without degradation. Moreover, RECAST enables automatic verification of constraint satisfaction via rule-based validators for quantitative constraints and LLM-based validators for qualitative ones, the verifiability provided by RECAST enables the design of reward functions for reinforcement learning, which further boosts model performance on complex and challenging tasks.
1 Introduction
Large language models (LLMs) have demonstrated remarkable capabilities in solving various NLP tasks (Ahn et al., 2024; Chakrabarty et al., 2024; Gómez-Rodríguez & Williams, 2023; Huang & Chang, 2023; Jiang et al., 2024a; Zhao et al., 2023; Zheng et al., 2023) and are widely used in practical scenarios (Gu et al., 2024; Liu et al., 2024; Sun, 2023; Thirunavukarasu et al., 2023). However, they still struggle with complex multi-constraint scenarios that are prevalent in real-world applications, such as constitutional AI systems requiring adherence to multiple principles simultaneously (Bai et al., 2022; Findeis et al., 2025; Kundu et al., 2023; Kyrychenko et al., 2025), and enterprise assistants managing detailed business rules (Michele et al., 2025; Grohs et al., 2023; Robino, 2025).
Recently, LLM-based autonomous agents have attracted increasing attention from both academia and industry. Surveys such as Wang et al. (2024b) highlight their tremendous potential in social sciences, natural sciences, and engineering, which fundamentally relies on the ability of LLMs to follow complex instructions. As shown in Figure 1, when instruction complexity increases, with multiple explicit constraints in a single prompt, even advanced models like GPT-4o (Achiam et al., 2023) show marked performance degradation, which has been emphasized in multiple benchmarks (Ferraz et al., 2024; Jiang et al., 2024b; Qin et al., 2024; Wen et al., 2024). Therefore, advancing the complex instruction-following capabilities of LLMs is crucial for promoting their practical deployment.
Although prior studies have explored methods to evaluate and improve the instruction-following abilities of LLMs, existing research remains limited in scope. Works such as Jiang et al. (2024b) and Wen et al. (2024) evaluated LLMs’ instruction-following performance and explored the use of LLMs themselves for evaluation. However, their datasets were manually constructed, lacked training sets, and thus fell short in terms of scalability for large-scale applications. While AutoIF (Dong et al., 2025) attempted to address data construction through an automated pipeline, it relied exclusively on code-verifiable constraints, which limits applicability to narrow domains and fails to capture broader, real-world use cases. More critically, most existing datasets impose only a small number of constraints (typically 3–5) and lack systematic evaluation benchmarks or training resources for complex instruction-following tasks. Previous automated approaches primarily rely on LLM-based instruction rewriting, which often yields homogenized constraints and limits scalability in generating large numbers of constraints. Consequently, they are inadequate for rigorously assessing and extending the capability boundaries of LLMs in realistic, complex application settings.
@@ 1 Introduction
To address these limitations and advance research in this area, we propose RECAST, a scalable data-synthesis framework that constructs instruction-following datasets of unprecedented complexity. Unlike prior approaches, our pipeline systematically mines instruction-following signals from existing data, thereby improving data utilization and fully exploiting the training value of available resources. Through the integration of diverse, verifiable constraints, RECAST enables more rigorous evaluation and more effective enhancement of LLMs’ ability to follow complex instructions.
Based on RECAST, we constructed a high-quality, diverse, and constraint-verifiable dataset, RECAST-30K, which encompasses a large number of constraints spanning diverse types. Using only 30K training samples, Llama3.1-8B-Base (Grattafiori et al., 2024) fine-tuned on RECAST-30K already outperforms other instruction-tuning datasets. Remarkably, it surpasses the corresponding instruct model—trained on a much larger-scale dataset—and even outperforms the substantially larger Llama-3.3-70B-Instruct in handling complex instruction-following scenarios.
Given that the constraints in RECAST-30K can be reliably verified using both rule-based and model-based validators, the dataset is inherently amenable to reinforcement learning. Building on this property, we design RLVC, which further enhances constraint compliance, yielding substantial improvements in multi-constraint satisfaction while preserving general capabilities. To summarize the highlights of our study:
  • We propose RECAST, an efficient, scalable, and low-cost data-synthesis framework that constructs complex instruction-following datasets of unprecedented complexity, aiming to challenge and enhance state-of-the-art models’ ability to follow complex instructions.
@@
  • We further exploit the verifiable constraints in RECAST-30K to design RLVC, a reinforcement-learning approach that leverages constraint-specific reward signals for simultaneous multi-objective optimisation, thereby fully exploiting the supervision potential of the dataset.
@@ 2 Realistic Extraction of Constraints for Augmented inStruction synThesis
We present RECAST, a comprehensive framework for improving LLMs’ ability to follow complex instructions. As shown in Figure 2, the RECAST pipeline operates through a three-stage process. First, we construct a rich constraint pool by extracting both rule-based and model-based constraints from seed data. Next, we synthesize enhanced instructions by selecting and naturally integrating appropriate constraints into original prompts. Finally, we generate consistent, constraint-compliant responses to these augmented instructions.
Figure 2: Overview of the RECAST framework and RLVC. Left: The RECAST pipeline generates complex instruction-following data through four steps: (1) seed data collection across diverse domains, (2) constraint construction with both rule-based and model-based verification methods, (3) instruction enhancement by integrating selected constraints, and (4) response synthesis ensuring constraints are satisfied. More detailed descriptions of the pipeline are provided in Appendix A. Top-right: Using RECAST-generated data to fine-tune LLMs through SFT. Bottom-right: RLVC framework leveraging constraint-specific verification to provide fine-grained rewards, guiding model optimization toward satisfying multiple constraints simultaneously.
@@ 2.1 Seed Data Collection
We select Tülu 3 Persona IF (Lambert et al., 2024) as the seed dataset for constructing RECAST-30K. This dataset spans diverse scenarios and task types, including code generation, creative writing, factual question answering, etc., rendering it an appropriate foundation for constraint extraction. A more detailed analysis of the characteristics of this seed dataset is provided in Appendix B.1.
@@ 2.2 Constraint Pool Construction
Rule-based Constraints Construction. Rule-based constraints represent objective requirements that can be verified through deterministic methods, including structural elements (paragraph count), lexical specifications (keyword inclusion/exclusion), and quantitative parameters (word limits). These constraints are particularly valuable as they provide unambiguous verification signals for model training and evaluation. To systematically extract these constraints, we implement nine specialized rule-based extractors (detailed in Appendix E.1) that analyze responses to identify verifiable properties. These extractors recognize specific syntactic patterns, keyword frequencies, numerical parameters, and structural characteristics that can be programmatically verified. This programmatic approach ensures that each rule-based constraint has a corresponding verification method, providing clear training signals for models to learn specific response characteristics.
Model-based Constraints Construction. Model-based constraints encompass subjective requirements that necessitate semantic understanding or qualitative judgment, such as stylistic elements (formality level), tonal qualities (politeness), and content characteristics (persuasiveness). These constraints are crucial for capturing the nuanced aspects of human communication that cannot be reduced to simple rules. Our construction process follows three key steps: First, we analyze each response from the seed datase to determine which constraint types from our taxonomy of 10 categories (detailed in Appendix E.2) are applicable, ensuring relevance to the specific content. Second, we employ LLMs to generate multiple concrete constraint instances for each applicable type, creating a comprehensive initial pool of model-based constraints. Third, we implement a filtering process to verify that each constraint is actually satisfied by its corresponding response, removing those that cannot be fulfilled. To validate the accuracy of filtering process, we conducted human evaluation on randomly selected constraint-response pairs, confirming that our methodology successfully identifies and retains only appropriate, verifiable constraints (detailed results in Appendix C.1).
@@ 2.3 Constraint-Augmented Instruction Synthesis
Constraint Selection. We employ LLMs to select a coherent subset of constraints from the pool that are relevant and applicable to each original instruction. This selection process ensures that only contextually appropriate constraints are considered for integration, avoiding the inclusion of irrelevant or contradictory requirements.
Instruction Enhancement. Once appropriate constraints are identified, we task multiple LLMs with integrating the selected constraints into the original instruction. Each model produces an enhanced instruction where constraints are incorporated naturally and coherently into the text.
Optimal Instruction Selection. To ensure quality and consistency, we implement a majority voting mechanism wherein multiple LLMs evaluate and rank the candidate integrated instructions. The evaluation criteria include linguistic fluency, semantic coherence, and constraint completeness. We select the highest-ranked instruction as the final constraint-augmented instruction. We also performed random sampling on optimal instruction selection for human evaluation, the results of human evaluation are shown in Appendix C.2.
@@ 2.4 Instruction-Consistent Response Synthesis
To address potential inconsistencies between constraint-augmented instructions and original responses in the seed dataset, we implement a two-phase process to ensure alignment while maintaining high response quality.
Diverse Response Generation. We generate multiple candidate responses to each constraint-augmented instruction using distinct LLMs. This multi-model approach introduces diversity into the candidate pool and mitigates potential biases inherent to any single model. Each model generates responses independently, ensuring a broad representation of possible approaches to satisfying the complex instruction.
Response Quality Assessment. We adopt a majority voting scheme considering constraint adherence, accuracy, conciseness, etc., to select the best response for each instruction. The chosen response is added to the dataset with its enhanced instruction. Human evaluation on sampled pairs (Appendix C.3) confirms the high quality of the selected responses.
This methodical approach to generation and selection ensures that our final dataset comprises constraint-rich and high-quality instruction–response pairs that demonstrate effective navigation of complex, multi-constraint scenarios. Each pair is annotated with its full set of constraints and corresponding validation method, facilitating downstream application and evaluation.
@@ 3 Reinforcement Learning with Verifiable Constraints as Rewards
Since the constraints in RECAST are equipped with corresponding verification methods, the dataset is inherently amenable to reinforcement learning. We therefore introduce RLVC, which leverages the verifiable nature of constraints to provide fine-grained reward signals during policy optimization. Building on SFT-trained models, we further apply RL to maximize the training value of RECAST-30K for enhancing complex instruction-following capabilities. This approach enables more targeted feedback on constraint satisfaction, thereby improving models’ ability to simultaneously address multiple complex requirements.
@@ 3.1 Constraint Verification Mechanisms
We implement a dual-mode verification scheme that enables precise evaluation of constraint satisfaction. Each constraint in an instruction is assessed through either rule-based or LLM-based verification methods. For rule-based constraints, we employ rule-based validators $V_{\text{rule-based}}(x,y,c_{i})$ that programmatically verify objective requirements through deterministic procedures. For model-based constraints that require evaluation of subjective qualities, we leverage LLM-based validators $V_{\text{model-based}}(x,y,c_{i})$. These validators assess more nuanced requirements that resist codification into explicit codes. We formalize the constraint verification function for an instruction $x$ with response $y$ and constraint $c_{i}$ as:
@@ 3.2 Verifiable Constraints as Reward Signals
The key innovation of RLVC lies in its exploitation of RECAST’s constraint verifiability feature. Traditional reinforcement learning approaches typically rely on a single, holistic reward signal that fails to identify which specific constraints are violated. In contrast, RECAST-30K provides individually verifiable constraints, enabling us to design a more informative reward mechanism. For an instruction $x$ containing multiple constraints $C=\{c_{1},c_{2},...,c_{n}\}$, we define the reward for a generated response $y$ as the average satisfaction rate across all constraints:
@@
This design provides an independent reward channel for each constraint, offering fine-grained feedback that guides the model toward simultaneously satisfying multiple constraints. Our RLVC framework combines verifiable constraints as reward signals with Group Relative Policy Optimization (GRPO) (Shao et al., 2024). Specifically, RLVC leverages the fine-grained feedback enabled by constraint verifiability and optimizes policies via GRPO’s comparative learning approach. This integration provides the model with clear guidance on improving constraint satisfaction across diverse instruction types. For completeness, we present the detailed formulation of GRPO, including its optimization objective, in Appendix D.5.